# Compress then Serve: Serving Thousands of LoRA Adapters with Little Overhead

### v1

@@ F.9 Reconstruction Error and Compression Rate
F.10 Reconstruction Error: Trained vs. Random
@@
F.11 Performance of Compressing Unseen/New LoRAs
Table 13 provides results for JD-full, where we project new unseen LoRAs into the shared basis. Our findings indicate that while compressing a new LoRA within the original space improves upon the baseline, it can degrades performance compared to uncompressed LoRAs
Model Type Method Type Tasks task036 task200 task220 task279 task281 task392 task520 task957 task1195 task1664 Base 11.11 21.02 26.81 8.59 48.08 12.78 15.06 46.42 61.35 43.01 Lora 87.37 94.00 99.00 94.33 80.25 94.00 99.00 47.94 97.86 100.00 10 16 14.49 27.85 82.15 62.23 53.49 63.31 43.66 48.73 69.81 54.85 32 14.64 46.41 95.83 69.89 53.63 72.66 74.86 48.12 71.25 57.27 64 19.63 67.00 98.16 71.66 55.32 80.51 79.60 48.13 76.02 63.92 50 16 16.09 82.00 97.16 70.14 55.26 76.01 83.74 47.71 77.92 61.46 32 44.28 83.00 99.00 72.66 56.68 86.55 86.29 47.74 79.44 68.21 64 57.76 88.00 99.00 74.00 56.86 92.33 88.05 47.75 84.29 71.29 100 16 22.18 84.00 98.00 71.66 55.49 82.38 88.12 48.27 78.71 66.81 32 48.45 86.00 99.00 73.66 56.85 88.32 88.05 48.02 81.81 67.19 64 60.60 90.00 99.00 74.00 59.63 92.33 90.00 48.18 84.92 71.52 500 16 60.99 86.00 99.00 73.00 56.58 90.66 91.00 47.20 85.16 71.21 32 64.09 90.00 99.00 73.66 59.58 92.00 91.00 47.37 86.84 75.23 64 66.29 89.00 100.00 71.33 59.79 93.00 92.00 47.56 88.29 74.98
Table 13: Compression new unseen LoRAs. ROUGE-L scores for various tasks and methods
F.12 Convergence
@@
where the tolerance threshold $\tau$ is set to 0.001. Due to the slow per-iteration computation times of the primary JD-Full algorithm, which quickly reaches an approximate optimum but then has a long tail of convergence for final digits of precision, we devised an alternative eigenvalue iteration algorithm (Appendix B.2) optimized for GPU acceleration. Our analysis indicates that adherence to this convergence criterion does not significantly alter the results.
@@
F.13 Out-of-distribution Performance (LoRA-hub)
For completeness, we incorporate results using the protocol of LoRA-hub [9]. That is, 100 LoRA-adapters are sampled, independent of the evaluation task, representing a measure of out-of-distribution performance. This also means that each result on a task is averaged across all 100 LoRA-adapters (as there is no a priori LoRA-to-task mapping). These results were obtained without normalizing the LoRA-adapters before applying the JD algorithms, a step we later identified as beneficial. We present performance comparison in Table 16. Table 15 presents the average agreement between uncompressed and compressed LoRA across 10 evaluation tasks. Results per task for JD-diagonal and JD-full are shown in Table 17 and Table 18, respectively.

### v2

@@ G.8 Reconstruction Error and Compression Rate
G.9 Reconstruction Error: Trained vs. Random
@@
G.10 Convergence
@@
where the tolerance threshold $\tau$ is set to 0.001. Due to the slow per-iteration computation times of the primary JD-Full algorithm, which quickly reaches an approximate optimum but then has a long tail of convergence for final digits of precision, we devised an alternative eigenvalue iteration algorithm (Appendix A.2) optimized for GPU acceleration. Our analysis indicates that adherence to this convergence criterion does not significantly alter the results.
@@
G.11 Out-of-distribution Performance (LoRA-hub)
For completeness, we incorporate results using the protocol of LoRA-hub (Huang et al., 2024). That is, 100 LoRA-adapters are sampled, independent of the evaluation task, representing a measure of out-of-distribution performance. This also means that each result on a task is averaged across all 100 LoRA-adapters (as there is no a priori LoRA-to-task mapping). These results were obtained without normalizing the LoRA-adapters before applying the JD algorithms, a step we later identified as beneficial. We present performance comparison in Table 14. Table 13 presents the average agreement between uncompressed and compressed LoRA across 10 evaluation tasks. Results per task for JD-diagonal and JD-full are shown in Table 15 and Table 16, respectively.